OpenStudy (anonymous):

Set up an equation and solve the problem. The perimeter of a rectangle is 50 inches, and the area is 144 square inches. Find the length and width of the rectangle.

OpenStudy (anonymous):

Let length =l breadth = b perimeter= 2(l+b)=50 area= lb=144 therefore substitue l=144/b in equation 1 and solve